# JL1-CD: A New Benchmark for Remote Sensing Change Detection and a Robust Multi-Teacher Knowledge Distillation Framework

### v2

@@ V-D Implementation Details
TTP changed +3.36 +5.69 -3.99 +3.39
V-E Experimental Results
Quantitative Comparison
Table III summarizes the numerical results of mIoU, mAcc, mPrecision, and mFscore for all methods on the JL1-CD test set, trained under the original, O-P, and MTKD strategies. As shown in the table, we observe that the O-P strategy does not lead to performance improvements for BAN-ViT-B, BAN-ViT-B-IN21k, FC-EF, FC-Siam-Conc, and FC-Siam-Diff. This suggests that the partition-based training method is not suitable for these algorithms, and thus cannot obtain sufficiently strong teacher models. Therefore, we do not continue with MTKD experiments for these methods. However, the O-P strategy or MTKD framework can improve the performance of all other algorithms to a certain extent. The following conclusions can be drawn from these results:
  1. 1)
    It is understandable that many models, such as SNUNet, BIT, and Changer-r18, perform better under the O-P strategy than under MTKD, as for well-trained models, the performance of teacher models should naturally surpass that of the student model. In the O-P strategy, the output from one of the teacher models is used as the final result, which allows its performance on the test set to exceed that of MTKD. However, it is equally noteworthy that several algorithms, such as STANet, IFN, and Changer-MiT-b1, perform worse under the O-P strategy than under MTKD, and in some cases, even worse than their original models. This is primarily because the original models fail to accurately estimate the CAR during the coarse detection stage on the test set. As a result, the subsequent selection of the wrong teacher model for further testing leads to a degradation in detection performance. This behavior highlights the potential shortcomings of the O-P strategy in certain scenarios. This finding also demonstrates that our MTKD method can enhance student models’ capabilities beyond those of the teacher models trained on partitioned data, without increasing computational or time costs during inference.
  2. 2)
    After MTKD optimization, the single Changer-MiT-b0 and Changer-MiT-b1 models can outperform the original TTP in terms of mIoU. Additionally, the TTP model, after MTKD optimization, shows improvements in mIoU and mFscore by 1.30% and 1.80%, respectively, setting a new SOTA.
  3. 3)
    The lightweight model TinyCD demonstrates competitive performance on the JL1-CD dataset, comparable to that of IFN (which has over 100 times more parameters). When trained within the MTKD framework, TinyCD’s mIoU and mFscore improve by 1.51% and 1.52%, respectively. This is crucial for resource-constrained edge device scenarios. However, while the O-P strategy significantly improves the performance of another lightweight model LightCDNet (with improvements of 3.49% in mIoU and 3.70% in mFscore), the MTKD framework leads to a performance degradation. This suggests that further improvements are needed in our approach when optimizing performance for lightweight models.
  4. 4)
    The O-P strategy shows the most significant improvement for Changer-MiT-s50 (with a 9.49% increase in mIoU and a 10.4% increase in mFscore), while the MTKD framework yields the most significant performance boost for HANet (with a 4.03% increase in mIoU and a 4.53% increase in mFscore).
Visual Comparison
In remote sensing CD, four main challenges are typically encountered: false alarms, missed detections, internal density, and boundary completeness. To visually assess the effectiveness of the O-P and MTKD strategies in addressing these challenges, we present the visual results of several algorithms in Fig. 7, where red indicates missed detections and blue indicates false alarms. In our experiments, we identify missed detections and false alarms as the most common issues, which are illustrated in the first six columns of Fig. 7. In Fig. 7(a), the original BAN algorithm exhibits many false alarms when detecting large CAR changes caused by PV panels and nearly fails to detect a newly constructed road on the right side of the image. After O-P optimization, false alarms related to PV panels are significantly reduced, but the road detection performance deteriorates. In contrast, MTKD achieves the best balance in detecting both small and large CAR changes, reducing false alarms related to PV panels while minimizing missed detections of the road. Fig. 7(c) similarly shows the results for detecting narrow roads. After MTKD optimization, SNUNet’s detection of the changed roads significantly improves in terms of completeness and connectivity, though some missed detections remain. his indicates that detecting changes caused by narrow roads remains a challenging task that requires further research. Our method not only improves the detection of large changes caused by buildings (Fig. 7(b)) and water body expansions (Fig. 7(e)) but also demonstrates notable advantages in detecting small object changes. Fig. 7(b) and Fig. 7(f) present results for small CAR scenarios. Whether qualitatively, as observed in the images, or quantitatively, as shown in the subsequent experiments, detecting small-scale changes remains a major challenge for all current CD algorithms. However, after optimization with O-P and MTKD, false alarms and missed detections in small-scale change scenarios are significantly improved. Of course, the improvement is limited; for instance, in the second and fourth columns, although MTKD detects more changes, it still cannot fully capture the tiny variations. This indicates that further improvements are needed in detecting small CAR changes. The last three columns illustrate the improvements in internal density and boundary completeness. We effectively reduce the occurrence of internal holes within the change areas, and the boundaries become more complete.
TABLE V: Impact of Different Numbers of Teacher Models on O-P and MTKD Performance
Method Strategy No. of $\mathcal{M}_{T}$ mIOU mAcc mPrecision mFscore
3 75.29 (+0.44) 81.40 (-0.44) 87.06 (+0.97) 81.32 (+0.34)
O-P 2 75.44 (+0.59) 81.96 (+0.12) 85.85 (-0.24) 81.51 (+0.53)
3 75.35 (+0.50) 81.76 (-0.08) 87.18 (+1.09) 81.28 (+0.30)
Changer (MiT-b0) MTKD 2 75.72 (+0.87) 82.30 (+0.46) 86.80 (+0.71) 81.66 (+0.68)
3 75.42 (-0.52) 81.67 (-0.32) 87.13 (-0.61) 81.43 (-0.50)
O-P 2 75.91 (-0.03) 82.11 (+0.12) 87.87 (+0.13) 81.97 (+0.04)
3 76.15 (+0.21) 82.85 (+0.86) 86.98 (-0.76) 82.13 (+0.20)
Changer (MiT-b1) MTKD 2 76.77 (+0.83) 83.38 (+1.39) 87.30 (-0.44) 82.66 (+0.73)
3 72.95 (-0.42) 79.71 (-0.60) 85.50 (+0.17) 79.12 (-0.53)
O-P 2 73.56 (+0.19) 80.76 (+0.45) 85.07 (-0.26) 79.92 (+0.27)
3 73.82 (+0.45) 80.32 (+0.01) 86.33 (+1.00) 79.91 (+0.26)
CGNet MTKD 2 73.78 (+0.41) 80.61 (+0.29) 85.67 (+0.34) 79.89 (+0.24)
3 76.69 (+1.64) 83.48 (+3.24) 87.27 (-2.55) 82.52 (+1.76)
O-P 2 76.65 (+1.60) 82.98 (+2.74) 87.39 (-2.43) 82.49 (+1.73)
3 76.85 (+1.80) 82.99 (+2.75) 88.05 (-1.77) 82.56 (+1.80)
TTP MTKD 2 76.31 (+1.26) 83.24 (+3.00) 86.81 (-3.01) 82.22 (+1.46)
@@ V-E Experimental Results
Figure 9: Visual comparison on the SYSU-CD dataset. Red denotes missed detections (FN). Blue indicates false alarms (FP). (a) Image at Time 1. (b) Image at Time 2. (c) Ground Truth. (d) Changer-MiT-b1 (Original). (e) Changer-MiT-b1 (MTKD). (f) CGNet (Original). (g) CGNet (MTKD). (h) TTP (Original). (i) TTP (MTKD).
Results on Different CAR Partitions
We select one model from each of the CNN, Transformer, and FM architectures and evaluate their performance across different CAR ranges. Fig. 8 summarizes the mIoU results of HANet, ChangeFormer-MiT-b1, and TTP on both the validation set (first row) and the test set (second row). The images are sorted by CAR in ascending order and divided into five equal partitions. Each bar in the figure represents the lower and upper bounds of CAR for each partition, while the line graph indicates the mIoU across the different partitions. From the figure, it can be observed that the detection performance of O-P and MTKD may actually decrease for images with high CAR (e.g., ChangeFormer-MiT-b1 and TTP). However, O-P and MTKD significantly enhance model performance on images with low CAR. In the first partition of the test set (CAR range 0.02% to 0.88%), O-P improves the mIoU for the three algorithms by 8.15%, 3.21%, and 6.55%, respectively, while MTKD improves it by 4.42%, 3.36%, and 2.85%. These results show that, due to the wide CAR range of the JL1-CD dataset, directly training existing algorithms using their original models can result in errors when detecting small changes. However, by partitioning small CAR regions in the O-P strategy and transferring the knowledge from the corresponding teacher model to the student model in MTKD, our methods achieve significant improvements in the accuracy of detecting tiny changes, which is consistent with the findings in Fig. 7.
Comparison of Results on Change and No-Change Classes
We further select all algorithms with an mIoU greater than 70% under the MTKD framework and compared their performance in detecting change and no-change regions. The experimental results are summarized in Table IV. Using the original models as a baseline, we analyze the contribution of MTKD to different metrics. Except for the ChangeFormer-MiT-b0 model, all MTKD-optimized models demonstrated significant improvements in both IoU and Fscore for the change regions. This indicates that the MTKD framework is more sensitive to the detection of change areas.
V-F Robustness of MTKD
Impact of Different Numbers of Teacher Models on MTKD Performance
In the previous subsection, we have demonstrated that the MTKD framework based on three teacher models can significantly improve the performance of various CD models. However, the 3-teacher configuration requires training five versions of the model for a single CD algorithm: $\mathcal{M}_{O}$, $\mathcal{M}_{T_{L}}$, $\mathcal{M}_{T_{M}}$, $\mathcal{M}_{T_{S}}$, and the student model $\mathcal{M}_{S}$. To alleviate the resource consumption associated with training these models and testing under the O-P strategy, we further explore the performance of the MTKD framework with only two teacher models. We set a threshold of $th=0.10$ to divide the dataset into small and large partitions and train the models $\mathcal{M}_{T_{S}}$ and $\mathcal{M}_{T_{L}}$ on these partitions, respectively. Subsequently, the student model $\mathcal{M}_{S}$ is trained using the MTKD framework. Except for the reduced number of teacher models, all other experimental settings remain unchanged. Using the original models as a baseline, Table V compares the metrics of models trained with the O-P strategy and the MTKD framework under different numbers of teacher models. The results indicate that two teacher models can still achieve significant performance improvements over the baseline. Notably, for Changer-MiT-b0 and Changer-MiT-b1, the two-teacher model setup yields even greater improvements compared to the three-teacher setup. For CGNet, the O-P strategy with two teachers achieves higher mIoU and mFscore, although the MTKD performance is lower in this case. For TTP, the performance of O-P and MTKD under the three teacher model still remains the strongest. These experiments confirm that the O-P strategy and MTKD framework remain effective with fewer teachers. In some cases, the 2-teacher configuration even outperforms the 3-teacher setup, indicating the flexibility of our approach.
Performance of MTKD on the SYSU-CD Dataset
We further evaluate the effectiveness of the MTKD framework on the SYSU-CD dataset. To ensure the representativeness and research significance of the experiments, we select one model from each of the three mainstream frameworks. The models are chosen with careful consideration:
  1. 1.
    TTP (FM-based) [34]: It represents the most recent and state-of-the-art FM-based model. Furthermore, it achieves the best performance on the JL1-CD dataset after optimization using our proposed MTKD framework.
  2. 2.
    Changer-MiT-b1 (Transformer-based) [14]: With only 13.355M parameters, this model achieves the second-highest mIoU and mFscore on the JL1-CD dataset, demonstrating its efficiency and effectiveness.
  3. 3.
    CGNet (CNN-based) [11]: Among all CNN-based models, CGNet has the largest parameter count and delivers the best performance, making it a strong representative of the CNN architecture.
We summarize the experimental results of these three algorithms on the SYSU-CD dataset, under different configurations of the number of teacher models, as well as their detection metrics for both change and no-change classes in Table VI. As shown in Table VI, regardless of the number of teacher models, the O-P strategy does not yield better results compared to the original models. The failure of the O-P strategy on the SYSU-CD dataset highlights its limitations. However, our MTKD approach still achieves notable success. After MTKD optimization, all models, whether with three or two teacher models, demonstrate higher mIoU and mFscore values. Notably, CGNet shows the most significant performance improvement in the three-teacher setup, while Changer-MiT-b1 and TTP perform better in the two-teacher setup. This contrasts with the test results on the JL1-CD dataset, suggesting that the optimal number of teacher models may vary depending on the dataset. This also indicates the need for further development of a method capable of automatically selecting the appropriate number of teachers based on the scenario. The Changer-MiT-b1 model with two teacher models shows the highest improvement in mIoU (+0.47%), while CGNet with three teacher models achieves the greatest improvement in mFscore (+0.60%).
In the last four columns of the table, we also present the detection metrics for both the change and no-change classes. For clarity, we show the absolute values obtained by the original models, while the results under the O-P strategy and the MTKD framework are represented as relative values, facilitating a more intuitive comparison of how our methods focus on the detection of change versus no-change areas. According to the MTKD setup with the optimal number of teachers, we observe higher improvements in the change class (e.g., the IoU and Fscore for the change class with CGNet in the three-teacher setup increased by 1.40% and 1.61%, respectively, while performance in the no-change class actually decreased). This demonstrates that MTKD is more sensitive to the detection of change areas, which is consistent with our findings on the JL1-CD dataset.
@@ V-F Robustness of MTKD
TABLE VI: Experimental Results on SYSU-CD Test Set (Best and Second-Best Results are in Bold and Underlined, Respectively)
@@
Method Strategy No. of $\mathcal{M}_{T}$ mIoU mAcc mPrecision mFscore Class IoU Acc Precision Fscore
Changer (MiT-b1) - - 75.49 84.25 86.38 82.38 unchanged 87.75 93.93 92.86 91.92
@@
O-P 3 75.32 83.88 85.86 82.19 unchanged +0.07 +0.19 +0.07 +0.12
changed -0.40 -0.93 -1.12 -0.52
2 75.47 83.79 86.34 82.31 unchanged +0.27 +0.48 -0.02 +0.27
changed -0.32 -1.40 -0.08 -0.41
MTKD 3 75.56 83.97 87.33 82.40 unchanged +0.15 +0.44 -0.12 +0.08
@@
O-P 3 75.97 84.54 86.20 82.66 unchanged -0.41 -0.79 +0.42 -0.26
changed +0.18 +0.71 -2.18 +0.15
2 75.83 83.71 86.69 82.83 unchanged -0.44 +0.24 -0.97 -0.38
changed -0.09 -1.99 +0.20 +0.60
MTKD 3 76.11 83.93 87.77 82.77 unchanged +0.05 +0.65 -0.73 +0.03
@@
The visual comparison is presented in Fig. 9, illustrating the detection results of these algorithms on four image pairs. Comparing the 4th and 5th columns, it is evident that MTKD significantly reduces the false alarm rate for Changer-MiT-b1. For example, in the 1st row, large areas of false detection are eliminated, and in the 3rd row, a small false-positive patch is removed. However, the 4th row exhibits an increase in false alarms. For CGNet, MTKD’s most notable contribution is the enhancement of internal compactness in detected regions. The original TTP model already produces satisfactory detection results, but MTKD further reduces false alarms. For instance, a large patch in the 1st row and a small patch in the 3rd row are effectively eliminated. However, the 1st row also shows an increase in missed detections. Overall, MTKD demonstrated substantial effectiveness on the SYSU-CD dataset, further validating its robustness.
VI Conclusion
In this work, we introduce a new benchmark dataset, JL1-CD, which significantly complements existing CD datasets by offering sub-meter resolution, a wide range of change types, and a large dataset scale. We also propose the MTKD framework, which significantly enhances the performance of CD models on dual-temporal remote sensing images with different change areas, without increasing time and computational complexity during inference. Our approach effectively improves the generalization and robustness of the model across different CD scenarios. Future work will focus on the adaptive selection of the number of teacher models and partition thresholds based on different scenarios, in order to enhance the flexibility and performance of the MTKD framework. Additionally, the evaluation of the MTKD framework on datasets with more complex and extreme change conditions, such as those affected by seasonal variations, severe weather events, or rapid urbanization, will be a critical direction for future research. This will help assess the robustness and adaptability of the framework under challenging real-world conditions.

### v3

@@ Appendix E Additional Experiments on the JL1-CD Dataset
TTP MTKD changed +3.36 +5.69 -3.99 +3.39
(a) FC-EF
(b) FC-Siam-Conc
(c) FC-Siam-Conc
(d) STANet
(e) IFN
(f) SNUNet
(g) BIT
(h) ChangeStar (FarSeg)
(i) ChangeStar (UPerNet)
(j) ChangeFormer (MiT-b0)
(k) ChangeFormer (MiT-b1)
(l) TinyCD
(m) HANet
(n) Changer (MiT-b0)
(o) Changer (MiT-b1)
(p) Changer (ResNet-18)
(q) Changer (ResNeSt-50)
(r) LightCDNet
(s) CGNet
(t) BAN
(u) TTP
Figure 8: mIoU of the benchmark CD methods across different CAR ranges. In each plot, the left y-axis represents CAR size, and the right y-axis represents mIoU.
@@
Figure 9: Visual comparison on the JL1-CD dataset. For each method, the images from left to right are: pre-event image, post-event image, ground truth, output of the original model, output of the O-P strategy, and output of the MTKD framework. Red indicates missed detections, while blue represents false alarms.
(a) FC-EF
(b) STANet
(c) TTP
Figure 10: Comparison between estimated CAR from the original model and ground truth CAR values.
Appendix F Failure Case Analysis
Figure 10 presents the classification matrix, CAR distribution comparison, and scatter plot between the coarse-predicted CAR of the original models and the GT CAR under the O-P strategy for FC-EF, STANet, and TTP. For FC-EF, the scatter plot reveals a large number of points where high-GT-CAR samples are mispredicted to have almost zero CAR, which indicates a severe issue of missed detections. This observation is further corroborated by the confusion matrix, where a substantial portion of “medium” and “large” regions are incorrectly classified as “small”. As a result, the predicted CAR distribution is heavily skewed, with the number of predicted “small” samples far exceeding that of the ground truth. In contrast, STANet exhibits the opposite trend: its original model tends to generate false alarms: approximately one-quarter of “small” samples are misclassified as “medium”, and more “medium” samples are misclassified as “large” than as “small”. Such misclassification of CAR regions by the original models leads to suboptimal performance under the O-P strategy, as reflected by the drop in mIoU of $7.03\%$ and $2.20\%$ for FC-EF and STANet, respectively.
On the other hand, the TTP algorithm, already exhibiting strong baseline performance, shows highly accurate CAR partitioning in its original model: the predicted CAR distribution closely matches that of the ground truth. Consequently, the O-P strategy provides a notable performance boost, improving mIoU and mFscore by $1.64\%$ and $1.76\%$, respectively. These analyses suggest that the effectiveness of the O-P strategy is highly dependent on the quality of the original model’s CAR predictions: its benefits are only realized when the baseline is sufficiently strong, while a poor baseline renders it less meaningful. Therefore, as a transitional approach, O-P has inherent limitations, whereas the MTKD method, which does not rely on coarse CAR estimation, is more robust and generally preferable.
Appendix G Additional Experiments on the SYSU-CD Dataset
The SYSU-CD dataset is a large-scale, relatively recent, and challenging dataset with a wide CAR range. Therefore, we select it as the second benchmark to evaluate the robustness of our proposed methods. Except for changing the patch size to 256 × 256, all other experimental settings are kept identical to those used for experiments on JL1-CD.
Table 8 presents additional quantitative results under different teacher model configurations, including detection metrics for both the change and no-change classes. As summarized in Table 8, our MTKD approach achieves consistent and notable improvements. After applying MTKD, all models, regardless of whether two or three teacher models are used, exhibit increased mIoU and mFscore values. Specifically, the FC-Siam-Diff model trained with two teacher models achieves the highest improvements in mIoU (+3.18%) and mFscore (+3.76%).
Among the evaluated models, SNUNet, ChangeFormer-MiT-b0, and CGNet demonstrate superior performance under the three-teacher setting, whereas FC-Siam-Diff, BiT, TinyCD, Changer-MiT-b1, and TTP perform better with two teachers. This observation suggests the need for future work on methods that can automatically select the optimal number of teachers depending on the scenario.
In the last four columns of Table 8, we report detection metrics separately for the change and no-change classes. The absolute values correspond to the original models, while the MTKD framework results are presented as relative improvements. We observe more substantial gains in detecting change areas. For instance, in the two-teacher setup, FC-Siam-Diff’s IoU and Fscore on the change class increase by 6.56% and 7.63%, respectively, whereas performance on the no-change class decreases slightly. These results suggest that MTKD enhances sensitivity toward change regions, consistent with our findings on the JL1-CD dataset.
Visual comparisons are shown in Figure 11. Missed detections remain the most common issue for the original models on this dataset. Both the O-P and MTKD strategies effectively mitigate these issues, with MTKD consistently achieving better results. Nonetheless, limitations remain. For example, while MTKD reduces missed detections in CGNet, some areas remain undetected. Similarly, in TinyCD, although both O-P and MTKD reduce missed detections, a few false positives and holes persist. These observations highlight promising directions for future research to further improve change detection performance.
@@ Appendix G Additional Experiments on the SYSU-CD Dataset
Table 8: Experimental results on SYSU-CD test set (best and second-best results are in bold and underlined, respectively).
@@
Method Strategy $\mathcal{M}_{T}$ mIoU mRec mPrec mFscore Class IoU Rec Prec Fscore
FC-Siam-Diff - - 62.04 70.78 85.80 69.10 unchanged 83.76 97.02 85.82 89.12
changed 40.32 44.54 85.78 49.09
[…]
changed +3.39 +4.70 -2.16 +3.61
2 65.22 75.14 82.83 72.86 unchanged -0.19 -2.60 +1.87 -0.12
changed +6.56 +11.33 -7.82 +7.63
[…]
changed 54.73 64.62 76.45 65.15
MTKD 3 70.61 79.85 83.15 78.29 unchanged +0.06 -0.43 +0.37 -0.06
changed +0.90 +1.72 -0.89 +1.04
[…]
changed +0.65 +0.79 +0.71 +0.78
BIT - - 67.16 76.03 84.27 74.76 unchanged 83.90 92.22 90.25 88.89
changed 50.42 59.85 78.30 60.62
[…]
changed +2.92 +5.32 -4.52 +3.12
[…]
changed +3.34 +5.87 -3.23 64.41
ChangeFormer (MiT-b0) - - 71.19 79.29 87.43 78.41 unchanged 86.33 94.32 91.03 90.82
changed 56.06 64.26 83.83 66.00
[…]
changed +1.60 +3.02 -2.12 +1.72
[…]
changed +1.15 +2.69 -3.91 +1.36
TinyCD - - 70.68 80.28 83.59 78.42 unchanged 85.14 93.96 89.91 89.92
changed 56.22 66.59 77.27 66.92
[…]
changed +0.53 +0.66 +0.28 +0.40
[…]
changed +0.69 +1.07 -0.48 +0.69
[…]
@@
[…]
@@
[…]
@@
Figure 11: Visual comparison on the SYSU-CD dataset. For each row, the images from left to right are: pre-event image, post-event image, ground truth, output of the original model, output of the O-P strategy with three teachers, output of the MTKD framework with three teachers, output of the O-P strategy with two teachers, and output of the MTKD framework with two teachers. Red indicates missed detections, while blue denotes false alarms.